OpenStudy (anonymous):

.Use the distributive property to rewrite the expression and solve. 6 94 = 6_____ + 6 4 = _____

OpenStudy (anonymous):

What is the first 6 suppose to be relative to the equation?

OpenStudy (anonymous):

i honestly have no idea

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

i suck at math

OpenStudy (shaadykhamlichi):

6*94=6*90+6*4=564

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

That makes sense

OpenStudy (shaadykhamlichi):

if you will write that don't write the stars

OpenStudy (shaadykhamlichi):

so i wish you will give me best response and be my fan!

OpenStudy (anonymous):

ok

OpenStudy (chosenmatt):

.